OpenStudy (anonymous):

Find the solutions to the following systems of equations. Which ordered pair "works"? -5x+y=-5 -4x+2y=2 a. (–8, –15) b. (–2, –15) c. (0, 1) d. (2, 5)

Vocaloid (vocaloid):

start by multiplying the first equation by (-2)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

nvm thats wrong hold on

OpenStudy (anonymous):

D

OpenStudy (anonymous):

would you like an explanation?

OpenStudy (anonymous):

@madison_taylor914 Yes please

OpenStudy (anonymous):

okay. well this is just a fill in the blank question. so at the top of the question you have 2 equations. they both have an x and y variable in them. then you have some plots. (x,y) so you just pick one of the answer choices and fill in the x and y in the equations. D is the answer because it is the only one that works in both equations. -5x+y=-5 -5(2) + 5 = -5 is a true statement. -4x+2y=2 -4(2) + 2(5)= 2 is also a true statement.

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

Yes! Thanks so much, @madison_taylor914.

OpenStudy (anonymous):

medal? if i helped

OpenStudy (anonymous):

thankyou